201982AM:

If the pressure after landing was 1.0 atm, what was the volume of the passenger's collapsed water bottle? (Assume the bottle collapsed easily and temperature didn't change.)

CheeseStick7793:

Using the ideal gas equation and Gay Lussac's law, the volume of the passenger's collapsed bottle after landing is 0.0075 L. What is ideal gas equation? The ideal gas equation shows the relationship between temperature, pressure, volume and number of moles. The volume of the passenger's collapsed bottle after landing is 0.0075 L. Given that; PV = nRT P = gas pressure V = volume of gas n = number of moles of gas R = molar gas constant T = temperature n = PV/RT Hence: P = 0.75 atm; V = 0.5 L; T = 295 K; R = 0.082 atmL/K.mol n = 0.75 * 0.5 / 0.082 * 295 n = 0.015 moles If the pressure after landing was 1.0 atm; V1n1 = V2n2 V2= V1n1/n2 V2 = 0.5 L * 0.015 moles/ 1 mole V2 = 0.0075 L The volume of the passenger's collapsed bottle after landing is 0.0075 L. Missing parts; A passenger on an airplane finished a 0.5 L bottle of water during the flight and replaced the lid. The cabin pressure and temperature were maintained at 0.75 atm and 295 K during the flight. After landing, he noticed his empty water bottle had collapsed. (12 points) i. How many moles of gas were in the water bottle?

CheeseStick7793:

hopefully this is useful to you

Rosee5656:

period